Three students S1, S2, and S3 are given a problem to solve. Consider the following events: U: At least one of S1, S2, and S3 can solve the problem, V: S1 can solve the problem, given that neither S2 nor S3 can solve the problem, W: S2 can solve the problem and S3 cannot solve the problem, T: S3 can solve the problem. For any event E, let P(E) denote the probability of E. If P(U) = 12, P(V) = 110, and P(W) = 112, then P(T) is equal to
- 1336
- 13
- 1960
- 14
Answer
(1) 1336
Solution
Let p1, p2, p3 be the (independent) probabilities that S1, S2, S3 solve the problem.
P(V) = p1 = 110 (independence makes the condition irrelevant).
P(U) = 1 - (1-p1)(1-p2)(1-p3) = 12 (1-p2)(1-p3) = 1/29/10 = 59.
(1-p2)(1-p3) = (1-p3) - p2(1-p3) = (1-p3) - 112.
1 - p3 = 59 + 112 = 2336 P(T) = p3 = 1336.
Answer: (A)